OpenStudy (anonymous):

find all real solutions of x^3−2x^2−3x+6=0

OpenStudy (turingtest):

x=2 is a solution so (x-2) is a factor. Divide synthetically to get a quadratic expression, which can be solved by either factoring or the quadratic formula.

OpenStudy (anonymous):

let f(x)=x^3−2x^2−3x+6 since f(2)=0 x-2 must be factor of f(x) use synthetic divsion or simple division you'll get than quadrtic equation. then solve that eqatin

OpenStudy (anonymous):

what told me that f(x) is 2

OpenStudy (turingtest):

nothing really. Trial and error. Descartes' Rule of signs says there are at most 2-positive roots, so I just plugged in x=1 and x=2 and got lucky.

OpenStudy (anonymous):

x=2 and\[\pm \sqrt{3}\]

OpenStudy (anonymous):

it's trial method, just put such value of x at which f(x)-0